# Digital Simulation of Non-Abelian Parafermions in Superconducting Circuits

### v3

@@ Abstract
Parafermions, which can be viewed as a fractionalized version of Majorana fermions, exhibit profound non-Abelian statistics and emerge in topologically ordered systems, while their realization in experiment has been challenging. Here we propose an experimental scheme for the digital simulation of parafermions and their non-Abelian braiding in superconducting (SC) circuits by realizing the $\mathbb{Z}_{d}$ plaquette model on a two-dimensional lattice. Two protocols using quantum circuits and non-destructive measurements are introduced to prepare the ground state, on which parafermion pairs are created by engineering dislocations. We then propose a generalized code deformation approach to realize the fusion and non-Abelian braiding statistics of parafermion modes, and show the application of this approach to fusing and braiding the $d=3$ parafermions. We also examine the experimental parameter regime to confirm the feasibility in SC devices. This work extends previous quantum simulation for twist defects in SC qubits to qudit systems, and may open up a way for parafermion-based high-dimensional topological quantum computing with experimental feasibility.
@@ I Introduction
The recent rapid advancement in quantum processors enables simulation of many-body quantum states and exploration of the intriguing physics they exhibit Aghaee et al. (2025); Chen et al. (2024); Xu et al. (2023); Andersen et al. (2023); Semeghini et al. (2021); De Léséleuc et al. (2019). This provides a way to study and manipulate exotic particles that have yet to be realized or are difficult to manipulate in real materials Nakamura et al. (2022, 2020); Rosenow et al. (2016); HANBURY BROWN and TWISS (1956); Bartolomei et al. (2020). A notable example is the simulation of anyons and topological defects in topologically ordered states (topological orders) Xu et al. (2023); Andersen et al. (2023); Kraus et al. (2013); Lu et al. (2009); Dai et al. (2017); Song et al. (2018); Fan et al. (2023); Xu et al. (2024). In particular, recent progress in digital simulation of twist defects in the $\mathbb{Z}_{2}$ toric code model with superconducting (SC) qubits demonstrates the fusion rules and (projective) non-Abelian braiding statistics Xu et al. (2023); Andersen et al. (2023). These efforts advance topologically protected quantum information processing, providing a foundation for the potential realization of topological quantum computing.
@@
Twist defects in the $\mathbb{Z}_{2}$ toric code model exhibit non-Abelian statistics of Majorana fermions, which is described by the Ising anyon model Bombin (2010); Kitaev (2006). Quantum computing schemes based on Majorana fermions have been widely studied, but their experimental realization in one- or two-dimensional condensed matter systems faces significant challenges Kitaev (2001); Ivanov (2001); Alicea et al. (2011); Alicea (2012); Aasen et al. (2016); Mourik et al. (2012); Deng et al. (2012); Rokhinson et al. (2012); Das et al. (2012); Wang et al. (2012); Churchill et al. (2013); Xu et al. (2014); Nadj-Perge et al. (2014); Chang et al. (2015); Albrecht et al. (2016); Wiedenmann et al. (2016); Bocquillon et al. (2017); Zhang et al. (2018); Wang et al. (2018); Fornieri et al. (2019); Ren et al. (2019); Jäck et al. (2019); Aghaee et al. (2023); Liu et al. (2017); Ahn et al. (2021); Pan and Sarma (2021); Yu et al. (2021); Takei et al. (2013); Knapp et al. (2020). More critically, braiding of Majorana fermions alone is insufficient for universal quantum computing Kitaev (2003a); Nayak et al. (2008). Universal quantum computing schemes based on Majorana fermions require additional non-topological assistance methods, making them not strictly topological Ahlbrecht et al. (2009); Ezawa (2024). $\mathbb{Z}_{d}$ parafermions generalize Majorana fermions (which are a special case for $d=2$) and exhibit more complex and nonlinear non-Abelian braiding statistics Benhemou et al. (2023); Levaillant et al. (2015). In contrast to Majorana fermions, parafermions can encode multi-level computational units, or qudits, which offer a larger state space for storing and processing information, as well as advantages in reducing circuit complexity and improving algorithm efficiency Wang et al. (2020). Braiding of parafermions allows the generation of the single-qudit Clifford group for any $d$, and the many-qudit Clifford group for odd $d$ Hutter and Loss (2016). Fault-tolerant non-Clifford gates can be performed by magic state distillation Anwar et al. (2012); Bravyi and Kitaev (2005); Campbell et al. (2012).
@@
Proposals for realizing parafermions typically involve systems with strong electron-electron interactions, which however pose serious experimental challenges Clarke et al. (2012); Cheng (2012); Clarke et al. (2013); Klinovaja et al. (2014); Klinovaja and Loss (2014a, b); Cobanera and Ortiz (2015); Alicea and Fendley (2016); Ebisu et al. (2017); Laubscher et al. (2019, 2020); Santos (2020); Schiller et al. (2020); Khanna et al. (2022); Teixeira and Dias da Silva (2022); Hong et al. (2024). On the other hand, to engineer the topological defects is an alternative general approach for realizing non-Abelian statistics within Abelian topological orders Barkeshli et al. (2013). Furthermore, the braiding of Abelian anyons facilitates performing non-Clifford gate in such phases. This provides a simpler way to explore parafermion defects in Abelian topologically ordered states. A typical class of stabilizer Hamiltonian models that realize Abelian topological orders is the $\mathbb{Z}_{d}$ plaquette models with qudits on a two-dimensional square lattice, in which certain lattice dislocations realize parafermions You and Wen (2012). As a stabilizer model, the digital simulation of its ground and excited states with a few quasiparticles is in principle achievable on experimental platforms with coherent control of qudits. The SC devices, such as transmons, which possess readily addressable higher energy levels, constitute promising candidates for implementing such a platform Blais et al. (2021). Recent works on SC devices have demonstrated high-fidelity single- and two-qutrit (qudit with $d=3$) entangling operations Cao et al. (2024); Roy et al. (2023); Luo et al. (2023); Yurtalan et al. (2020); Morvan et al. (2021); Goss et al. (2022, 2024). Nevertheless, realization of the parafermion modes and implementation of their non-Abelian statistics necessitate qualitatively more complicated quantum gate operators and involve addressing non-trivial physical challenges in comparison with the Majorana fermions Andersen et al. (2023); Xu et al. (2023), and it remains unclear how to achieve such quantum simulation in the real SC circuits. More precisely, while the generalization from Majorana fermions to parafermion modes may appear straightforward at the level of mathematical model construction, its physical realization remains significantly non-trivial. First, the fundamental Pauli operators and two-qudit gates are unitary but not Hermitian, requiring careful consideration of both the operations and their complex conjugates in circuit implementations. Second, the multi-level structure of qudits necessitates novel physical mechanisms to control transitions between different levels and to implement entangling operations involving higher levels. Examples include the two-photon process for controlling the 00-$2$ transition in qutrits Yurtalan et al. (2020) to realize high-fidelity generalized Hadamard gate, and implementing control gates based on tunable cross-Kerr interactions Luo et al. (2023); Goss et al. (2022). Third, parafermion exhibit richer non-Abelian braiding statistics and fusion rules compared to Majorana fermions, making it physically non-trivial to identify suitable observables for characterizing their braiding and fusion properties. This work aims to integrate recently developed qudit techniques in SC devices to develop an experimentally feasible scheme for simulating parafermions and their fusion and braiding, providing insights into the simulation of high-spin quantum many-body systems and high-dimensional quantum computing.
@@
Here, we summarize the main results of our work. In this work, we propose an experimentally feasible scheme to simulate parafermions and their non-Abelian braiding in SC circuits, focusing on the $\mathbb{Z}_{d}$ plaquette model on a planar surface. We introduce two efficient ground state preparation protocols: one using quantum circuits and the other employing non-destructive measurements. By engineering dislocations in the square lattice, pairs of parafermions are created. With a novel generalized code deformation approach, as proposed in the present work, these parafermions are moved on the planar surface, enabling a feasible way for their fusion and braiding operations. The concrete example of $d=3$ parafermions is studied in detail. The real parameter conditions are examined, with the feasibility of our scheme in experiment being confirmed. We note that while the present simulation is designed for the SC devices, it is also compatible with other experimental platforms capable of implementing the gates that we use. Our work provides a theoretical foundation for the imminent digital simulation of parafermions in quantum computing platforms, and shall advance topologically protected high-dimensional quantum information processing.
@@ II The Model
Figure 1: Illustration of the model with a pair of prafermions. (a) The $\mathbb{Z}_{d}$ plaquette model on an checkerboard lattice with a dislocation colored by light blue, where the small white dots represent qudits. The red cross marks the defect associated with the dislocation, identified as the parafermion mode $\gamma$. The red dotted line indicates the branch cut connecting the two parafermions. The arrowed dark blue curves represent string operators that create $e-\overline{e}$ and $m-\overline{m}$ excitations, which reside on the light and dark plaquettes, respectively. In our convention, the $e$ or $m$ particle is located at the start of the string operator. Stabilizers $O_{p}$ on both dark and light plaquettes are shown. Graphical representation of the generalized Pauli operators, commutation relation, and stabilizers are shown. (b) An illustration of the energy diagram and microwave driving of a typical transmon as a multilevel system. The frequency difference $\omega_{12}-\omega_{01}$ characterizes the anharmonicity. (c) Dislocations are engineered by replacing two neighboring stabilizers with a single one. Applying a two-qudit quantum circuit moves and separates the parafermion defects. (d) Schematic of the cross-Kerr interaction that is used to realize two-qutrit gates. (e) A specific configuration for realizing the tunable cross-Kerr interaction using a frequency-tunable coupler between two transmon qutrits.
We start by considering the $\mathbb{Z}_{d}$ plaquette model on the two-dimensional square lattice with a checkerboard pattern [Fig. 1(a)]. On each vertex $i$, there is a qudit with $d$ basis states $\left|{m_{i}}\right\rangle$ ($m_{i}=0,1,2...,d-1$). Physically, the qudit can be realized by the intrinsic multi-dimensional Hilbert space of SC devices, such as transmons. Their coherence properties and sufficient anharmonicity enable transmons to function as qudits, with the lowest $d$ energy levels serving as the qudit basis states [Fig. 1(b)]. The Hilbert space is spanned by the qudit states on all the vertices. The Hamiltonian is
@@
where the labels of the qudits are shown in the top of Fig. 1(a), and the sum of plaquettes $p$ is for all the light and dark plaquettes. Here, $Z$, $X$, $Z^{\dagger}$, and $X^{\dagger}$ are generalized Pauli operators for qudits. They are defined by the properties: $X^{d}=\mathbbm{1},\,Z^{d}=\mathbbm{1},\,XZ=\omega ZX,$ where $\omega=e^{\frac{2\pi i}{d}}$ and act on the qudit basis as $Z\left|{m}\right\rangle=\omega^{m}\left|{m}\right\rangle$ and $X\left|{m}\right\rangle=\left|{m-1\,(\text{mod}\quad d)}\right\rangle$. The experimental realization of single-qudit operations can be achieved through microwave driving resonant with specific qudit state subspaces. For $d=3$ (i.e., a qutrit), any single-qutrit gate can be decomposed into a sequence of operations acting on two-level subspaces Blok et al. (2021); Morvan et al. (2021). Accounting for the AC Stark effect due to anharmonicity implements high-fidelity single-qutrit gates Blok et al. (2021); Liu et al. (2016). It is convenient to introduce a graphical representation [Fig. 1(a)] for the generalized Pauli operators and their commutation relations. With this representation, the stabilizer $O_{p}$ is described by a loop operator around the plaquette $p$, which satisfies $O_{p}^{d}=1$ and has eigenvalues $1$, $\omega$,…, $\omega^{d-1}$. Stabilizers on different plaquettes commute with each other; therefore, the Hamiltonian Eq. (1) is exactly solvable. The ground state $\left|{\Psi_{0}}\right\rangle$ satisfies:
@@
and the ground-state projector is
@@
The model defined by Eq. (1) provides a Hamiltonian realization of the topological gauge theory with gauge group $\mathbb{Z}_{d}$. The stabilizers on light (dark) plaquettes correspond to local Gauss’s law (flatness condition) constraints. The ground state, defined by Eq. (2), is the state where all local constraints are satisfied, while excited states arise from violations of these constraints. For $d=2$, the stabilizer has only two eigenvalues $1$ and $-1$, indicating whether a local constraint is satisfied or violated. The elementary excitations—violations of Gauss’s law and flatness condition—are termed charge ($e$) and flux ($m$) excitations, respectively, following its physical meaning in gauge theory. In the $\mathbb{Z}_{2}$ gauge theory, these excitations are self-dual, satisfying $e^{2}=1$ and $m^{2}=1$. For $d\geq 3$, however, charge and flux excitations are no longer self-dual and can appear in multiple powers. Specifically, the elementary excitations of Eq. (1) include charge excitations $e,e^{2},...,e^{d-1}$ on light plaquettes and flux excitations $m,m^{2},...,m^{d-1}$ on dark plaquettes [see Appendix. A for an algebraic characterization of these excitations]. More general excitations involve composites $e^{k}m^{d-k}$, known as dyons. When a charge excitation $e$ encircles a flux excitation $m$ counterclockwise, or vice versa, the system acquires a phase factor $\overline{\omega}\equiv\omega^{d-1}$ due to the Aharonov-Bohm effect. Consequently, charge, flux, and dyon excitations are Abelian anyons, created by string operators (graphically represented in Fig. 1(a)). The ground state of Eq. (1) exhibits the Abelian topological order characterized by the quantum double $D(\mathbb{Z}_{d})$ Kitaev (2003b).
@@
To incorporate non-Abelian statistics with the Abelian topological order model, we need to engineer dislocations. The dislocation in the square lattice that we consider is an extended object formed by tilting the edges along a line [the light blue region in Fig. 1(a)]. Associated with the dislocation are two point-like defects at the ends, characterized by trivalent vertices on the pentagon plaquettes. In the $\mathbb{Z}_{d}$ plaquette model, the dislocation is associated with deforming stabilizers $O_{p}$ and $O_{p^{\prime}}$ on two neighbouring plaquettes to a single stabilizer $O_{p^{\prime}}O_{p}$ [Fig. 1(c)]. This deformation effectively removes the edge between the plaquettes and reduces the number of stabilizers by one. Therefore, $N$ such dislocations result in the $Nd$-fold ground-state degeneracy, with each defect having a quantum dimension $\sqrt{d}$ that indicates the non-Abelian nature of the defect. These non-Abelian defects manifest parafermionic behavior, which arises from the physics of defect-induced electric-magnetic exchange symmetry. To illustrate this, we create an $e-\overline{e}$ charge pair by applying a $Z$ gate to a qudit, and move the $\overline{e}$ particle through the branch cut connecting two defects using a string operator [Fig. 1(a)]. The $\overline{e}$ excitation transforms into the $\overline{m}$ excitation, indicating an $e-m$ exchange. This shows that the defect identifies the dyon exictation $e^{k}m^{d-k}$ ($k=1,2,...,d-1$) with the trivial topological charge $1$, giving rise to the following fusion rules:
@@
where $\gamma$ denotes the defects. Equation (4) is exactly the fusion rules of parafermions. For clarity, we henceforth refer to these defects as parafermions. By applying certain two-qudit gates, whose general forms are shown later, we can move the parafermions apart in space. This allows to realize the fusion and braiding of parafermions. For convenience, we focus on the qutrit case, i.e., the $\mathbb{Z}_{3}$ plaquette model in the following. Generalizing the methods to larger $d$ is straightforward.

### v4

@@ Abstract
Parafermions, which can be viewed as a fractionalized version of Majorana modes, exhibit profound non-Abelian statistics and emerge in topologically ordered systems, while their realization in experiment has been challenging. Here we propose an experimental scheme for the digital simulation of parafermions and their non-Abelian braiding statistics in superconducting (SC) circuits by realizing the $\mathbb{Z}_{d}$ plaquette model on a two-dimensional lattice. Two protocols using quantum circuits and non-destructive measurements are introduced to prepare the ground state, on which parafermion pairs are created by engineering dislocations. We then propose a generalized code deformation approach to realize the fusion and non-Abelian braiding of parafermion modes, and show the application of this approach to the $\mathbb{Z}_{3}$ parafermions. We also examine the experimental parameter regime to confirm the feasibility of our scheme in SC devices. This work extends previous quantum simulation of twist defects in SC qubits to qudit systems, and opens up a way for parafermion-based high-dimensional topological quantum computing with experimental feasibility.
@@ I Introduction
The recent rapid advancement in quantum processors enables simulation of many-body quantum states and exploration of the intriguing physics they exhibit [1, 2, 3, 4, 5, 6]. This provides a way to study and manipulate exotic particles that have yet to be realized or are difficult to manipulate in real materials [7, 8, 9, 10, 11]. A notable example is the simulation of anyons and topological defects in topologically ordered states (topological orders) [3, 4, 12, 13, 14, 15, 16, 17]. In particular, recent progress in digital simulation of twist defects in the $\mathbb{Z}_{2}$ toric code model with superconducting (SC) qubits demonstrates the (projective) non-Abelian braiding statistics [3, 4]. These efforts advance topologically protected quantum information processing, providing a foundation for the potential realization of topological quantum computing.
@@
Twist defects in the $\mathbb{Z}_{2}$ toric code model exhibit non-Abelian statistics of Majorana fermion modes, which is described by the Ising anyon model [18, 19]. Quantum computing schemes based on Majorana modes have been widely studied, but their experimental realization in one- or two-dimensional condensed matter systems faces significant challenges [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. More critically, braiding of Majorana modes alone is insufficient for universal quantum computing [49, 50]. Universal quantum computing schemes based on Majorana modes require additional non-topological assistance methods, making them not strictly topological [51, 52]. $\mathbb{Z}_{d}$ parafermions generalize Majorana modes (which are a special case for $d=2$) and exhibit more complex and nonlinear non-Abelian braiding statistics [53, 54]. In contrast to Majorana modes, parafermions can encode multi-level computational units, or qudits, which offer a larger state space for storing and processing information, as well as advantages in reducing circuit complexity and improving algorithm efficiency [55]. Braiding of parafermions allows the generation of the single-qudit Clifford group for any $d$, and the many-qudit Clifford group for odd $d$ [56]. Fault-tolerant non-Clifford gates can be performed by magic state distillation with high fidelity [57, 58, 59].
@@
Proposals for realizing parafermions typically involve systems with strong electron-electron interactions, which however pose serious experimental challenges [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. On the other hand, to engineer the topological defects is an alternative general approach for realizing non-Abelian statistics within Abelian topological orders [76]. This provides a simpler way to realize parafermion defects in Abelian topologically ordered states. Furthermore, the braiding of Abelian anyons of the underlying state facilitates performing topological non-Clifford gates. A typical class of stabilizer Hamiltonian models that realize Abelian topological orders is the $\mathbb{Z}_{d}$ plaquette models with qudits on a two-dimensional square lattice, in which certain lattice dislocations realize parafermions [77]. As a stabilizer model, the digital simulation of its ground state and excited states with a few quasiparticles is in principle achievable on experimental platforms with coherent control of qudits. The SC devices, such as transmons, which possess readily addressable higher energy levels, constitute promising candidates for implementing such platforms [78]. Recent works on SC devices have demonstrated high-fidelity single- and two-qutrit (qudit with $d=3$) entangling operations [79, 80, 81, 82, 83, 84, 85]. Nevertheless, realization of the parafermion modes and implementation of their non-Abelian statistics necessitate qualitatively more complicated quantum gate operators and involve addressing non-trivial physical challenges in comparison with the Majorana modes [4, 3], and it remains unclear how to achieve such quantum simulation in the real SC circuits. More precisely, while the generalization from Majorana mdoes to parafermion modes may appear straightforward at the level of mathematical model construction, its physical realization remains significantly non-trivial. First, the fundamental single- and two-qudit gates are unitary but not Hermitian, requiring careful consideration of both the operations and their Hermitian conjugates in circuit implementations. Second, the multi-level structure of qudits necessitates novel physical processes to control level transitions and to implement entangling operations involving higher levels. Examples include the two-photon process for controlling the 00-$2$ transition in qutrits [82] to realize high-fidelity generalized Hadamard gate, and implementing control gates based on tunable cross-Kerr interactions [81, 84]. Third, parafermions exhibit richer non-Abelian braiding statistics and fusion rules compared to Majorana modes, making it physically non-trivial to identify suitable observables for characterizing their braiding and fusion properties. This work aims to address these challenges and integrate recently developed qudit techniques in SC devices to develop an experimentally feasible scheme for simulating parafermions and their fusion and braiding, providing insights into the parafermion-based high-dimensional topological quantum computing.
@@
Here, we summarize the main results of our work. In this work, we propose an experimentally feasible scheme to simulate parafermions and their fusion and non-Abelian braiding in SC circuits, focusing on the $\mathbb{Z}_{d}$ plaquette model on a planar surface. We introduce two efficient ground state preparation protocols: one using quantum circuits and the other employing non-destructive measurements. By engineering dislocations in the lattice, pairs of parafermions are created. With a novel generalized code deformation approach, as proposed in the present work, these parafermions are moved on the planar surface, enabling a feasible way for their fusion and braiding operations. The concrete example of $d=3$ parafermions is studied in detail. The real parameter conditions are examined, with the feasibility of our scheme in experiment being confirmed. We note that while the present simulation is designed for the SC devices, it is also compatible with other experimental platforms capable of implementing the gates that we use. Our work provides a theoretical foundation for the imminent digital simulation of parafermions in quantum computing platforms, and shall advance topologically protected high-dimensional quantum information processing.
@@ II The Model
Figure 1: Illustration of the model with a pair of prafermions. (a) The $\mathbb{Z}_{d}$ plaquette model on a checkerboard lattice with a dislocation colored by light blue, where the small white dots represent qudits. The red cross marks the defect associated with the dislocation, identified as the parafermion mode $\gamma$. The red dotted line indicates the branch cut connecting the two parafermions. The arrowed dark blue curves represent string operators that create $e-\overline{e}$ and $m-\overline{m}$ excitations, which reside on the light and dark plaquettes, respectively. In our convention, the $e$ or $m$ particle is located at the start of the string operator and its dual is located at the end of the string operator. Stabilizers $O_{p}$ on both dark and light plaquettes are shown. Graphical representations of the generalized Pauli operators, commutation relation, and stabilizers are shown. (b) An illustration of the energy diagram and microwave driving of a typical transmon as a multilevel system. The frequency difference $\omega_{12}-\omega_{01}$ characterizes the anharmonicity. (c) Dislocations are engineered with two neighboring stabilizers being replaced with a single one. Applying a two-qudit quantum circuit moves and separates the parafermion defects. (d) Schematic of the cross-Kerr interaction that is used to realize two-qutrit gates. (e) A configuration that realizes the tunable cross-Kerr interaction using a frequency-tunable coupler, as demonstrated in Ref. [81].
We start by considering the $\mathbb{Z}_{d}$ plaquette model on the two-dimensional square lattice with a checkerboard pattern [Fig. 1(a)]. On each vertex $i$, there is a qudit with $d$ basis states $\left|{m_{i}}\right\rangle$ ($m_{i}=0,1,2...,d-1$). Physically, the qudit can be realized by the intrinsic multi-dimensional Hilbert space of SC devices, such as transmons. Their coherence properties and sufficient anharmonicity enable to function as qudits, with the lowest $d$ energy levels serving as the qudit basis states [Fig. 1(b)]. The Hilbert space is spanned by the qudit states on all the vertices. The Hamiltonian is:
@@
where the labels of the qudits are shown in the top of Fig. 1(a), and the sum of plaquettes $p$ is for all the light and dark plaquettes. Here, $Z$, $X$, $Z^{\dagger}$, and $X^{\dagger}$ are generalized Pauli operators for qudits. They are defined by the properties: $X^{d}=\mathbbm{1},\,Z^{d}=\mathbbm{1},\,XZ=\omega ZX,$ where $\omega=e^{\frac{2\pi i}{d}}$ and act on the qudit basis as $Z\left|{m}\right\rangle=\omega^{m}\left|{m}\right\rangle$ and $X\left|{m}\right\rangle=\left|{m-1\,(\text{mod}\quad d)}\right\rangle$. The experimental realization of single-qudit operations can be achieved through microwave driving resonant with specific qudit state subspaces. For $d=3$ (i.e., a qutrit), any single-qutrit gate can be decomposed into a sequence of operations acting on two-level subspaces [86, 83], and implemented with high fidelity by accounting for the AC Stark effect arising from anharmonicity [86, 87]. It is convenient to introduce a graphical representation [Fig. 1(a)] for the generalized Pauli operators and their commutation relations. With this representation, the stabilizer $O_{p}$ is described by a loop operator around the plaquette $p$, which satisfies $O_{p}^{d}=1$ and has eigenvalues $1$, $\omega$,…, $\omega^{d-1}$. Stabilizers on different plaquettes commute with each other; therefore, the Hamiltonian Eq. (1) is exactly solvable. The ground state $\left|{\Psi_{0}}\right\rangle$ satisfies:
@@
and the ground-state projector is:
@@
The model defined by Eq. (1) provides a Hamiltonian realization of the topological gauge theory with gauge group $\mathbb{Z}_{d}$. The stabilizers on light (dark) plaquettes correspond to local Gauss’s law (flatness condition) constraints. The ground state, defined by Eq. (2), is the state where all local constraints are satisfied, while excited states arise from violations of these constraints. For $d=2$, the stabilizer has only two eigenvalues $1$ and $-1$, indicating whether a local constraint is satisfied or violated. The elementary excitations—violations of Gauss’s law and flatness condition—are termed charge ($e$) and flux ($m$) excitations, respectively, following its physical meaning in gauge theory. In the $\mathbb{Z}_{2}$ gauge theory, these excitations are self-dual, satisfying $e^{2}=1$ and $m^{2}=1$. For $d\geq 3$, however, charge and flux excitations are no longer self-dual and can appear in multiple powers. Specifically, the elementary excitations of Eq. (1) include charge excitations $e,e^{2},...,e^{d-1}$ on light plaquettes and flux excitations $m,m^{2},...,m^{d-1}$ on dark plaquettes [see Appendix. A for an algebraic characterization of these excitations]. More general excitations involve composites $e^{k}m^{d-k}$, known as dyons. When a charge excitation $e$ encircles a flux excitation $m$ counterclockwise, or vice versa, the system acquires a phase factor $\overline{\omega}\equiv\omega^{d-1}$ due to the Aharonov-Bohm effect. Consequently, charge, flux, and dyon excitations are Abelian anyons, created by string operators [graphically represented in Fig. 1(a)]. The ground state of Eq. (1) exhibits the Abelian topological order characterized by the quantum double $D(\mathbb{Z}_{d})$ [88].
@@
To incorporate non-Abelian statistics with the Abelian topological order model, we need to engineer dislocations. The dislocation in the square lattice that we consider is an extended object formed by tilting the edges along a line [the light blue region in Fig. 1(a)]. Associated with the dislocation are two point-like defects characterized by trivalent vertices on the pentagon plaquettes. In the $\mathbb{Z}_{d}$ plaquette model, the dislocation is associated with deforming stabilizers $O_{p}$ and $O_{p^{\prime}}$ on two neighbouring plaquettes to a single stabilizer $O_{p^{\prime}}O_{p}$ [Fig. 1(c)]. This deformation effectively removes the edge between the plaquettes and reduces the number of stabilizers by one. Therefore, $N$ such dislocations result in the $Nd$-fold degeneracy, with each defect having a quantum dimension $\sqrt{d}$ that indicates the non-Abelian nature of the defect. These non-Abelian defects manifest parafermionic behavior, whose underlying physics is the defect-induced electric-magnetic exchange symmetry. To illustrate this, consider creating an $e-\overline{e}$ charge pair by applying a $Z$ gate to a qudit, and moving the $\overline{e}$ particle through the branch cut connecting two defects using a string operator [Fig. 1(a)]. The $\overline{e}$ excitation transforms into the $\overline{m}$ excitation, indicating an $e-m$ exchange. This shows that the defect identifies the dyon exictation $e^{k}m^{d-k}$ ($k=1,2,...,d-1$) with the trivial topological charge $1$, giving rise to the following fusion rules:
@@
where $\gamma$ denotes the defects. Equation (4) is exactly the fusion rules of parafermions. For clarity, we henceforth refer to these defects as parafermions. By applying certain two-qudit operations, whose circuit realizations are shown later, we can move the parafermions apart in space. This allows to realize the fusion and braiding of parafermions. For convenience, we focus on the qutrit case, i.e., the $\mathbb{Z}_{3}$ plaquette model in the following. Generalizing the methods to larger $d$ is straightforward.